The intensity of sound heard by Jonathan when his headphones is “on” his ear is around \ (3.6x10-4 W/m2 \). Assuming the initial distance is around 4.6cmwhen the headphone is on his ear, what is the intensity when he removes the headset, and is a distance 7.4cm from his ear?

Note: Use this notation in writing units with exponents that are more than one: s= s2



1
Expert's answer
2021-12-07T09:55:21-0500
"I=\frac{4.6^2}{7.4^2}3.6\cdot10^{-4}=1.4\cdot10^{-4}\frac{W}{m^2}"